Hugo Enterprises (HE) is a small oil services firm in the Houston area. For the upcoming production period, Hugo must forecast demand for X200. Historical demand is given in the table below. Use exponential smoothing with alpha = 0.7 to forecast period 4. (Round all calculations to two decimal places)

Period

Demand

1

500

2

536

3

600

Forecast for Period 4 =                    Units
Answer #1

F(t) = F(t-1) + (Alpha * (A(t-1) - F(t-1)))

FORECAST 2 = 500 + (0.7 * (500 - 500)) = 500

FORECAST 3 = 500 + (0.7 * (536 - 500)) = 525.2

FORECAST 4 = 525.2 + (0.7 * (600 - 525.2)) = 577.56